# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.2.2 Uniform estimates
$\quad\leq C\big{(}\|\mu^{\gamma,k}\|_{L^{q}}+\|\varphi^{\gamma,k}\|_{L^{q}}+\|\sigma^{\gamma,k}\|_{L^{q}}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{q}}\big{)},\quad\text{for}\ q\in[2,6].$ (3.47)
Recalling the estimates [30, (3.15), (3.16)], we infer from (3.42) that
$\left|\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\right|$ $\leq C\big{(}\|\nabla\mu^{\gamma,k}\|+\|\sigma^{\gamma,k}-\overline{\sigma^{\gamma,k}}\|+\|\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big{)}\|\nabla\varphi^{\gamma,k}\|+C$
$\leq C\big{(}1+\|\nabla\mu^{\gamma,k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big{)}.$ (3.48)
The estimates (3.42), (3.47), (3.48) combined with the Sobolev embedding theorem and the Poincaré–Wirtinger inequality yield
$\quad\leq C\left(1+\|\nabla\mu^{\gamma,k}\|+\left|\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\right|+\|\sigma^{\gamma,k}\|_{H^{1}}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{q}}\right)$
$\quad\leq C\big{(}1+\|\nabla\mu^{\gamma,k}\|+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{q}}\big{)},\quad\text{for}\ q\in[2,6].$ (3.49)
On the other hand, from Hölder’s and Young’s inequalities, the Sobolev embedding theorem, (3.42) and (3.44), we find
$\|\sigma^{\gamma,k}\|_{L^{4}}^{4}$ $\leq\|\sigma^{\gamma,k}\|_{L^{12}}^{2}\|\sigma^{\gamma,k}\|_{L^{3}}\|\sigma^{\gamma,k}\|$
$\leq\|\sigma^{\gamma,k}\|_{L^{12}}^{4}+C\|\sigma^{\gamma,k}\|_{L^{3}}^{2}$
$\leq C_{3}\big{(}1+\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}\big{)},$ (3.50)
where $C_{3}>0$ depends on coefficients of the system, norms of the initial data and $\Omega$, but is independent of $\gamma$, $k$ and $T$. Combining (3.46), (3.49) and (3.50), we infer that
$\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{4}}^{4}+C_{4}\|\sigma^{\gamma,k}\|_{L^{4}}^{4}+\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}$
$\quad\leq C\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)},$ (3.51)
where $0<C_{4}<\frac{1}{C_{3}}$. It follows from (3.42), (3.51) and Gronwall’s lemma that for all $t\in[0,T]$, it holds
$\|\sigma^{\gamma,k}(t)\|_{L^{4}}^{4}$ $\leq\|\sigma^{\gamma,k}(0)\|_{L^{4}}^{4}e^{-C_{4}t}+C$
$\quad+C\int_{0}^{t}\big{(}\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)}\,\mathrm{d}s$
$\leq C+C\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s,$ (3.52)
and for all $t\in[0,T-1]$,
$\int_{t}^{t+1}\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}\,\mathrm{d}s$ $\leq\|\sigma^{\gamma,k}(0)\|_{L^{4}}^{4}e^{-C_{4}t}+C$
$\quad+C\int_{0}^{t+1}\big{(}\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)}\,\mathrm{d}s$
$\leq C+C\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s.$
Here, the positive constant $C$ is independent of $\gamma$, $k$ and $T$.
Testing (3.36c) by $(\sigma^{\gamma,k})^{5}$ yields
$\frac{1}{6}\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{6}}^{6}+\frac{5}{9}\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}=-\int_{\Omega}\chi\Delta\varphi^{\gamma,k}(\sigma^{\gamma,k})^{5}\,\mathrm{d}x.$ (3.53)
@@
From the Sobolev embedding theorem and Poincaré’s inequality, it follows that
@@
$\|\sigma^{\gamma,k}\|_{L^{18}}^{3}=\|(\sigma^{\gamma,k})^{3}\|_{L^{6}}\leq C\left(\|\nabla(\sigma^{\gamma,k})^{3}\|+\left|\int_{\Omega}(\sigma^{\gamma,k})^{3}\,\mathrm{d}x\right|\right).$ (3.54)
Then by (3.42), (3.54), (3.52) and Young’s inequality, we get
$\|(\sigma^{\gamma,k})^{5}\|_{L^{\frac{6}{5}}}^{2}=\|\sigma^{\gamma,k}\|_{L^{6}}^{10}$ $\leq\|\sigma^{\gamma,k}\|_{L^{4}}^{\frac{40}{7}}\|\sigma^{\gamma,k}\|_{L^{18}}^{\frac{30}{7}}\leq\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}+C.$
The term on the right-had side of (3.53) can be estimated by
$-\int_{\Omega}\chi\Delta\varphi^{\gamma,k}(\sigma^{\gamma,k})^{5}\,\mathrm{d}x$ $\leq C|\chi|\|\Delta\varphi^{\gamma,k}\|_{L^{6}}\|(\sigma^{\gamma,k})^{5}\|_{L^{\frac{6}{5}}}$
$\leq C\chi^{2}\|\Delta\varphi^{\gamma,k}\|^{2}_{L^{6}}+\frac{2}{9}\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}+C.$ (3.55)
Hence, it follows from (3.53) and (3.55) that
$\frac{1}{6}\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{6}}^{6}+\frac{1}{3}\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}\leq C\|\Delta\varphi^{\gamma,k}\|^{2}_{L^{6}}+C.$ (3.56)
From Hölder’s and Young’s inequalities, the Sobolev embedding theorem, (3.42), (3.52) and (3.54), we find
$\|\sigma^{\gamma,k}\|_{L^{6}}^{6}$ $\leq\|\sigma^{\gamma,k}\|_{L^{18}}^{3}\|\sigma^{\gamma,k}\|_{L^{3}}^{2}\|\sigma^{\gamma,k}\|_{L^{6}}$
$\leq\|\sigma^{\gamma,k}\|_{L^{18}}^{6}+C\|\sigma^{\gamma,k}\|_{L^{6}}^{2}$
$\leq C_{3}^{\prime}\big{(}1+\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}\big{)},$ (3.57)
Combining (3.49), (3.56) and (3.57), we infer that
$\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{6}}^{6}+C_{4}^{\prime}\|\sigma^{\gamma,k}\|_{L^{6}}^{6}+\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}$
$\quad\leq C\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)},$
where $0<C_{4}^{\prime}<\frac{1}{C_{3}^{\prime}}$. Hence, for all $t\in[0,T]$, it holds
$\|\sigma^{\gamma,k}(t)\|_{L^{6}}^{6}$ $\leq\|\sigma^{\gamma,k}(0)\|_{L^{6}}^{6}e^{-C_{4}^{\prime}t}+C$
$\quad+C\int_{0}^{t}\big{(}\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)}\,\mathrm{d}s$
$\leq C+C\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s,$ (3.58)
and for all $t\in[0,T-1]$,
$\int_{t}^{t+1}\|\nabla(\sigma^{\gamma,k})^{3}\|^{2}\,\mathrm{d}s$ $\leq\|\sigma^{\gamma,k}(0)\|_{L^{6}}^{6}e^{-C_{4}^{\prime}t}+C$
$\quad+C\int_{0}^{t+1}\big{(}\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)}\,\mathrm{d}s$
$\leq C+C\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s.$ (3.59)
The positive constant $C$ in (3.58) and (3.59) is independent of $\gamma$, $k$ and $T$.
Fourth estimate. Now we estimate the time derivative of $\sigma^{\gamma,k}$. It follows from (3.36c) and (3.52) that
$\|\partial_{t}\sigma^{\gamma,k}\|_{(H^{1})^{\prime}}^{2}$ $\leq\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\|\bm{v}^{k}\|_{\bm{L}^{4}}^{2}\|\sigma^{\gamma,k}\|_{L^{4}}^{2}$
$\leq\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\|\bm{v}^{k}\|_{\bm{L}^{4}}^{2}\left(C+C\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right),$ (3.60)
which together with (3.42) and Poincaré’s inequality yields
$\int_{0}^{T}\|\partial_{t}\sigma^{\gamma,k}(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s\leq C+C\left(\int_{0}^{T}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right)^{2}.$ (3.61)
Next, testing (3.36c) by $-\Delta\sigma^{\gamma,k}$, we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\sigma^{\gamma,k}\|^{2}+\|\Delta\sigma^{\gamma,k}\|^{2}$ $=\int_{\Omega}\bm{v}^{k}\cdot\nabla\sigma^{\gamma,k}\Delta\sigma^{\gamma,k}\,\mathrm{d}x+\chi\int_{\Omega}\Delta\varphi^{\gamma,k}\Delta\sigma^{\gamma,k}\,\mathrm{d}x$
$\leq\|\bm{v}^{k}\|_{\bm{L}^{\infty}}\|\nabla\sigma^{\gamma,k}\|\|\Delta\sigma^{\gamma,k}\|+|\chi|\|\Delta\varphi^{\gamma,k}\|\|\Delta\sigma^{\gamma,k}\|$
$\leq\frac{1}{2}\|\Delta\sigma^{\gamma,k}\|^{2}+\|\bm{v}^{k}\|_{\bm{L}^{\infty}}^{2}\|\nabla\sigma^{\gamma,k}\|^{2}+\chi^{2}\|\Delta\varphi^{\gamma,k}\|^{2},$ (3.62)
which together with (3.49) yields
$\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\sigma^{\gamma,k}\|^{2}+\|\Delta\sigma^{\gamma,k}\|^{2}$
$\quad\leq 2\|\bm{v}^{k}\|_{\bm{L}^{\infty}}^{2}\|\nabla\sigma^{\gamma,k}\|^{2}+C\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+\gamma^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}\big{)}.$
By Gronwall’s lemma and (3.42), we get
$\|\nabla\sigma^{\gamma,k}(t)\|^{2}\leq\big{[}\|\nabla\sigma^{\gamma,k}(0)\|^{2}+C(1+T)\big{]}\exp\left(2\int_{0}^{T}\|\bm{v}^{k}(s)\|_{\bm{L}^{\infty}}^{2}\,\mathrm{d}s\right),\quad\forall\,t\in[0,T],$ (3.63)
and
$\int_{0}^{T}\|\Delta\sigma^{\gamma,k}(s)\|^{2}\,\mathrm{d}s$ $\leq\|\nabla\sigma^{\gamma,k}(0)\|^{2}+C(1+T)$
$\quad+2\|\bm{v}^{k}\|_{L^{\infty}(0,T;\bm{L}^{\infty}(\Omega))}^{2}\int_{0}^{T}\|\nabla\sigma^{\gamma,k}(s)\|^{2}\,\mathrm{d}s.$ (3.64)
From
$\|\partial_{t}\sigma^{\gamma,k}\|\leq\|\bm{v}^{k}\|_{\bm{L}^{\infty}}\|\nabla\sigma^{\gamma,k}\|+\|\Delta\sigma^{\gamma,k}\|+|\chi|\|\Delta\varphi^{\gamma,k}\|,$
we can further deduce that
$\int_{0}^{T}\|\partial_{t}\sigma^{\gamma,k}(s)\|^{2}\,\mathrm{d}s\leq C,$ (3.65)
where $C$ is independent of $\gamma$ by depends on $k$ and $T$ (via $\|\bm{v}^{k}\|_{L^{\infty}(0,T;\bm{L}^{\infty}(\Omega))}$).
@@
Fifth estimate. Testing (3.36a) by $\partial_{t}\mu^{\gamma,k}$ and integrating over $\Omega$, after integration by parts we obtain
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\mu^{\gamma,k}\|^{2}+\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x+\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\mu^{k}\,\mathrm{d}x+\alpha(\overline{\varphi^{\gamma,k}}-c_{0})\int_{\Omega}\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x=0.$
A direct computation yields that
@@
$\quad-\chi\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\sigma^{\gamma,k}\,\mathrm{d}x+\beta\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\mathcal{N}\big{(}\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big{)}\,\mathrm{d}x.$
Next, by exploiting the incompressibility and the no-slip boundary condition of the velocity field $\bm{v}^{k}$, we have
$\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x$
$\quad=\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\big{(}\gamma\partial_{t}^{2}\varphi^{\gamma,k}-\Delta\partial_{t}\varphi^{\gamma,k}+\varPsi_{0}^{\prime\prime}(\varphi^{\gamma,k})\partial_{t}\varphi^{\gamma,k}-\theta_{0}\partial_{t}\varphi^{\gamma,k}-\chi\partial_{t}\sigma^{\gamma,k}\big{)}\,\mathrm{d}x$
$\qquad+\beta\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\mathcal{N}\big{(}\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big{)}\,\mathrm{d}x$
$\quad=\frac{\mathrm{d}}{\mathrm{d}t}\left(\gamma\int_{\Omega}\bm{v}^{\gamma,k}\cdot\nabla\varphi^{\gamma,k}\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x\right)+\gamma\int_{\Omega}\partial_{t}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varphi^{\gamma,k}\,\mathrm{d}x$
@@
$\qquad-\int_{\Omega}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varPsi_{0}^{\prime}(\varphi^{\gamma,k})\,\mathrm{d}x+\theta_{0}\int_{\Omega}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varphi^{\gamma,k}\,\mathrm{d}x$
$\qquad-\chi\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\partial_{t}\sigma^{\gamma,k}\,\mathrm{d}x-\beta\int_{\Omega}\bm{v}^{k}\cdot\nabla\mathcal{N}\big{(}\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big{)}\varphi^{\gamma,k}\,\mathrm{d}x.$
Besides, it follows from (3.39) that
@@
Collecting the above identities, we arrive at
$\frac{\mathrm{d}}{\mathrm{d}t}\left[\frac{1}{2}\|\nabla\mu^{\gamma,k}\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x+\alpha\big{(}\overline{\varphi^{\gamma,k}}-c_{0}\big{)}\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\right]$
@@
$\quad=\theta_{0}\|\partial_{t}\varphi^{\gamma,k}\|^{2}-\gamma\int_{\Omega}\partial_{t}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varphi^{\gamma,k}\,\mathrm{d}x-\int_{\Omega}\big{(}\nabla^{T}\bm{v}^{k}\nabla\varphi^{\gamma,k}\big{)}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x$
$\qquad-\int_{\Omega}\big{(}\nabla^{2}\varphi^{\gamma,k}\bm{v}^{k}\big{)}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x+\int_{\Omega}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varPsi^{\prime}_{0}(\varphi^{\gamma,k})\,\mathrm{d}x$
$\qquad-\theta_{0}\int_{\Omega}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k}\varphi^{\gamma,k}\,\mathrm{d}x+\chi\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\sigma^{\gamma,k}\,\mathrm{d}x$
@@
$\qquad+\chi\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\partial_{t}\sigma^{\gamma,k}\,\mathrm{d}x+\beta\int_{\Omega}\bm{v}^{k}\cdot\nabla\mathcal{N}\big{(}\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big{)}\varphi^{\gamma,k}\,\mathrm{d}x$
@@
The first six terms $I_{1},\cdots,I_{6}$ on the right-hand side of (3.66) can be handled as in [2, Section 2] with suitable modifications. In view of (3.39), (3.42) and Poincaré’s inequality, we have

### v2

@@ 3.2.2 Decoupled auxiliary problem for $(\varphi,\mu)$
$\quad\leq C\big(\|\mu^{\gamma,k}\|_{L^{q}}+\|\varphi^{\gamma,k}\|_{L^{q}}+\|\sigma^{k}\|_{L^{q}}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{q}}\big),\quad\forall\,q\in[2,6],$ (3.73)
@@
where $C>0$ depends on $\Omega$ and $q$. From (3.67), (3.70), (3.72), (3.73), the Sobolev embedding theorem and the Poincaré–Wirtinger inequality, we obtain
$\|\varphi^{\gamma,k}(t)\|_{W^{2,q}}+\|\varPsi_{0}^{\prime}(\varphi^{\gamma,k}(t))\|_{L^{q}}$
@@
$\quad\leq C\left(\|\nabla\mu^{\gamma,k}(t)\|+\left|\int_{\Omega}\mu^{\gamma,k}(t)\,\mathrm{d}x\right|+\|\sigma^{k}(t)\|_{L^{q}}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|_{L^{q}}+1\right)$
$\quad\leq C\big(\|\nabla\mu^{\gamma,k}(t)\|+\|\sigma^{k}(t)\|_{L^{q}}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|_{L^{q}}+1\big),\quad\forall\,q\in[2,6],$ (3.74)
for almost all $t\in[0,T-1]$. The constant $C>0$ is independent of $\gamma$, $k$ and time. Taking $q=2$, from (3.72) and (3.74), we find
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\left(\|\varphi^{\gamma,k}(s)\|_{H^{2}}^{2}+\|\varPsi_{0}^{\prime}(\varphi^{\gamma,k}(s))\|^{2}\right)\,\mathrm{d}s\leq C.$
@@
Fourth estimate. Testing (3.66a) by $\partial_{t}\mu^{\gamma,k}$ and integrating over $\Omega$, after integration by parts, we get
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\mu^{\gamma,k}\|^{2}+\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x+\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x+\alpha(\overline{\varphi^{\gamma,k}}-c_{0})\int_{\Omega}\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x=0.$
By the definition of $\mu^{\gamma,k}$, a direct computation yields
@@
$\quad-\chi\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\sigma^{k}\,\mathrm{d}x+\beta\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\mathcal{N}\big(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big)\,\mathrm{d}x.$
Next, from the construction of $\bm{v}^{k}$, $\sigma^{k}$ and $(\varphi^{\gamma,k},\mu^{\gamma,k})$, using integration by parts, we observe that
$\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\mu^{\gamma,k}\,\mathrm{d}x$
$\quad=\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\big(\gamma\partial_{t}^{2}\varphi^{\gamma,k}-\Delta\partial_{t}\varphi^{\gamma,k}+\varPsi_{0}^{\prime\prime}(\varphi^{\gamma,k})\partial_{t}\varphi^{\gamma,k}-\theta_{0}\partial_{t}\varphi^{\gamma,k}-\chi\partial_{t}\sigma^{k}\big)\,\mathrm{d}x$
$\qquad+\beta\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\mathcal{N}\big(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big)\,\mathrm{d}x$
$\quad=\frac{\mathrm{d}}{\mathrm{d}t}\left(\gamma\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x\right)+\gamma\int_{\Omega}(\partial_{t}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x$
@@
$\qquad-\int_{\Omega}(\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varPsi_{0}^{\prime}(\varphi^{\gamma,k})\,\mathrm{d}x+\theta_{0}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x$
$\qquad-\chi\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\sigma^{k}\,\mathrm{d}x-\beta\int_{\Omega}\big[\bm{v}^{k}\cdot\nabla\mathcal{N}\big(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big)\big]\varphi^{\gamma,k}\,\mathrm{d}x.$
Finally, using (3.68), we can get
@@
Combining the above identities, we arrive at the following differential equality:
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\|\nabla\mu^{\gamma,k}\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x+\alpha\big(\overline{\varphi^{\gamma,k}}-c_{0}\big)\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\right)$
@@
$\quad=\theta_{0}\|\partial_{t}\varphi^{\gamma,k}\|^{2}-\gamma\int_{\Omega}(\partial_{t}\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x-\int_{\Omega}\big(\nabla^{T}\bm{v}^{k}\nabla\varphi^{\gamma,k}\big)\cdot\nabla\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x$
$\qquad-\int_{\Omega}\big(\nabla^{2}\varphi^{\gamma,k}\bm{v}^{k}\big)\cdot\nabla\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x+\int_{\Omega}(\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varPsi^{\prime}_{0}(\varphi^{\gamma,k})\,\mathrm{d}x$
$\qquad-\theta_{0}\int_{\Omega}(\bm{v}^{k}\cdot\nabla\partial_{t}\varphi^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x+\chi\int_{\Omega}\partial_{t}\varphi^{\gamma,k}\partial_{t}\sigma^{k}\,\mathrm{d}x$
@@
$\qquad+\chi\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\sigma^{k}\,\mathrm{d}x+\beta\int_{\Omega}\big[\bm{v}^{k}\cdot\nabla\mathcal{N}\big(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big)\big]\varphi^{\gamma,k}\,\mathrm{d}x$
@@
The first six terms $I_{1},\cdots,I_{6}$ on the right-hand side of (3.75) can be estimated as in [3, Section 2] with suitable modifications. To this end, using (3.68), (3.72) and the Poincaré–Wirtinger inequality, we infer that